Prove 0! =1 with proper logic and maths.
Answer #1


Broof of 0! = 1 n! - nx (n-1) x (n-2) (m-1)! = (n-1) x (n-2) but n=1 n! = nx (n-1)! I! x (1-1) 1 IXO! | = 0! CS Scanned with